# Toroidal Cartesian Products Where One Factor is $3$-Connected

### v1

@@
Abstract.
In this paper, we provide a partial classification of Cartesian products of graphs that embed on the torus where one factor is $3$-connected. We show that if $G\boxempty H$ embeds on the torus and $G$ is $3$-connected, then $H$ must be a path on two vertices, $P_{2}$, and we provide significant evidence that $G$ must be outer-cylindrical. At the same time, we show that if $G$ is outer-cylindrical, then $G\boxempty P_{2}$ embeds on the torus. This all leads us to conjecture that if $G$ is $3$-connected then $G\boxempty H$ embeds on the torus if and only if $G$ is outer-cylindrical and $H=P_{2}$. We also determine the genera of $K_{4}\boxempty P_{3}$ and $H_{i}\boxempty P_{2}$ where $H_{i}$ is a planar, $3$-connected, non-outer-cylindrical graph derived from a wheel graph $W_{i}$ via two vertex splits and one edge addition.
1. Introduction
@@
Classifying embeddings of graphs on surfaces is a well-studied topic in topological graph theory. Here, we are interested in embeddings of a Cartesian product of two graphs $G$ and $H$, denoted $G\boxempty H$, which is defined to have vertex set $V(G)\times V(H)$ and a pair of vertices $(u,v),(u^{\prime},v^{\prime})\in V(G)\times V(H)$ determine an edge in $G\boxempty H$ if and only $u=u^{\prime}$ and $vv)\in E(H)$ or $uu^{\prime}\in E(G)$ and $v=v^{\prime}$. We refer the reader to [7] for further background on graph Cartesian products. The fiber structures of graph Cartesian products can sometimes be leveraged to efficiently build embeddings and determine genera of such graphs; see [10], [11], [13], [14], [15], [17], [20], [21] for some examples. In particular, it would be interesting to know if one can determine the set of all Cartesian products of graphs that embed on a fixed surface. Such a classification was established for planar graphs, that is, graphs that embed in the plane, in the work of Behzad and Mahmodian [4]. More recently, the second author along with Abell and McDermott have provided a complete classification of Cartesian products of graphs that embed in the projective plane in [1].
@@
In this paper, we are interested in classifying which Cartesian product of graphs are toroidal, i.e., embed on the torus. While the classifications of planar and projective planar Cartesian products were reasonably straightforward, we expect the classification of toroidal Cartesian products to be far more complicated. To start, the work of Kuratowski [8] and Wagner [19] showed there are only two minor-minimal obstructions to planarity, namely $K_{5}$ and $K_{3,3}$, and Archdeacon in [3] showed that there is a complete list of 35 graphs that are minor-minimal obstructions to embedding on the projective plane. However, the list of minor-minimal obstructions for embedding on the torus isn’t completely known, and this list contains at least 17 thousand forbidden minors; see [12]. In addition, when restricting to Cartesian products of graphs, we see far more flexibility in toroidal embeddings than planar or projective planar embeddings with regards to (vertex) connectivity. If $G\boxempty H$ is planar or projective planar, then both $G$ and $H$ are each at most $2$-connected. This follows from the fact that $K_{4}\boxempty P_{2}$ is both non-planar and non-projective planar, and any $3$-connected graph contains a $K_{4}$ minor; see Lemma 4.2. However, in Section 5, we provide infinite classes of $3$-connected and $4$-connected graphs $\{G_{i}\}$ such that each $G_{i}\boxempty P_{2}$ is toroidal. Thus, even a partial classification of toroidal Cartesian products could be a far more complicated task than the planar and projective planar cases.
Here, we provide some major first steps towards this classification problem, with the additional assumption that one factor is $3$-connected. A simple combinatorial argument shows that if $G\boxempty H$ is toroidal, then $G$ and $H$ can each be at most $4$-connected in Proposition 4.1. Then in Theorem 4.3 we show that if $G$ is $3$-connected, then any toroidal $G\boxempty H$ must have $H=P_{2}$. This proof follows from the fact that $K_{4}\boxempty P_{3}$ is not toroidal, which is proved in Theorem 3.4. In Section 5, we provide two explicit constructions of infinite classes of toroidal $G_{i}\boxempty P_{2}$ and observe that any such $G_{i}$ in these constructions is planar and outer-cylindrical; see the beginning of Section 5 for the definition of outer-cylindrical. This all leads us to Conjecture 5.4, which states that if $G$ is $3$-connected, then $G\boxempty H$ is toroidal if and only if $H=P_{2}$ and and $G$ outer-cylindrical. Our work proves the backwards direction of this conjecture and provides strong evidence in favor of the forward direction. Theorem 4.3 already shows that $H=P_{2}$ holds under the hypotheses of the forward direction. Section 6 provides further evidence that the forward direction of this conjecture holds by showing the existence of an infinite set of planar, $3$-connected, non-outer-cylindrical graphs $\{H_{i}\}$ such that each $H_{i}\boxempty P_{2}$ does not embed on the torus and $H_{i}$ is “not far” in some sense from a wheel graph, $W_{i}$, which is an outer-cylindrical $3$-connected graph.
@@
There are two significant by-products of our work on this classification problem. In Section 3 we prove that the genus of $K_{4}\boxempty P_{3}$ is two, highlighted in Theorem 3.4. This work requires a careful combinatorial analysis of all potential embeddings of this graph on a torus, which employs techniques similar to [5]. In particular, even though the girth of $K_{4}\boxempty P_{3}$ is three, one cannot hope for a triangular embedding since every edge in a $P_{3}$ fiber does not lie on a $3$-cycle. Thus, determining the genera of such a graph is not a straightforward task. In addition, in Section 6, we prove Theorem 6.10, which shows that $H\boxempty P_{2}$ has genus two, where $H$ is a specified planar, $3$-connected, non-outer-cylindrical graph. Parts of this proof follow the structure of the proof of Theorem 3.4, but the analysis is much more complex. As far as the authors can tell, both Theorem 3.4 and Theorem 6.10 provide novel contributions to the topological graph theory literature.
We would like to acknowledge the Furman University Mathematics Department for financially supporting this work via the Summer Mathematics Undergraduate Research Fellowships.
2. Background
@@
In this paper, we assume all graphs are simple, connected, and non-trivial (in the sense of being nonempty and not a single vertex), unless otherwise noted. We use $S_{g}$ to denote an orientable, closed, connected surface of genus $g\in\mathbb{N}\cup\{0\}$. Given an embedding of a graph $G$ on a surface $S_{g}$ via a function $f:|G|\rightarrow S_{g}$ where $|G|$ represents a geometric realization of $G$, we sometimes use $f(G)$ to denote this embedding or just $G$ when the embedding is clear from context. If each component of $S_{g}\setminus f(G)$ is homeomorphic to an open disk, then we say this embedding is a $2$-cell embedding. The genus of a graph $G$, denoted $\gamma(G)$, is the minimum $g\in\mathbb{N}\cup\{0\}$ such that $G$ embeds on $S_{g}$. An embedding of $G$ that realizes the minimal genus is called a minimal embedding. We will mainly be interested in $2$-cell embeddings of graphs, and since the work of Youngs [22] shows that each minimal embedding of $G$ is a $2$-cell embedding, we usually don’t need to make a distinction when determining the genera of a graph. We refer the reader to [6] for further background on graph embeddings on surfaces.
@@
Fact 1: Each vertex of $G$ has an open $2$-cell neighborhood in $S$.
Fact 2: Each edge of $G$ borders precisely two (not necessarily unique) faces.
@@
Facts (1) and (2) are general facts about $2$-cell embeddings. Fact (3) follows from (2) and since any edge in a $P_{n}$ fiber is not contained in a $3$-cycle. Fact (4) follows from (2). This collection of facts will play an important role in ruling out potential toroidal embeddings of graphs in Section 3 and Section 6.
A graph $G^{\prime}$ is a minor of $G$ if $G^{\prime}$ can be obtained by a series of vertex deletions, edge deletions, or edge contractions on $G$. Another way to find a lower bound for the genus of a graph $G$ is by finding the genus (or lower bound on the genus) for a subgraph or minor $G$. Since we will be interested in Cartesian products of graphs in this paper, we first confirm that the product of minors from each factor results in a minor of the product.
Proposition 2.2.
@@
Let $G$ and $H$ be graphs and suppose $G^{\prime}$ be a minor of $G$. Recall that the vertex sets and edge sets of $G\boxempty H$ are defined as $V(G\boxempty H)=V(G)\times V(H)$ and $E(G\boxempty H)=(V(G)\times E(H))\cup(E(G)\times V(H))$. It suffices to show $G^{\prime}\boxempty H$ is a minor of $G\boxempty H$ under an edge deletion, vertex deletion, or edge contraction, which we now handle as three separate cases.
Case 1: Let $G^{\prime}=G-e$ for some $e\in E(G)$. Then $G^{\prime}\boxempty H$ can be obtained from $G\boxempty H$ by deleting edge $(e,v)\in E(G)\times V(H)$, for each $v\in V(H)$.
@@
Case 3: Given edge $e\in E(G)$ with endpoints $a$ and $b$, we edge contract $e$ in $G$ to create $G^{\prime}$ which now has a new vertex $u\in V(G^{\prime})$ to which $a$ and $b$ have been identified. Then $G^{\prime}\boxempty H$ can be obtained from $G\boxempty H$ by contracting edge $(e,v)\in E(G)\times V(H)$, for each $v\in V(H)$. ∎
Corollary 2.3.
@@
3. The genus of $K_{4}\boxempty P_{3}$
In this section, we will show that $\gamma(K_{4}\boxempty P_{3})=2$. Since every $3$-connected graph contains a $K_{4}$ minor (see Lemma 4.2), this will play a pivotal role in classifying toroidal Cartesian products where one of the factors is $3$-connected.
@@ 3. The genus of $K_{4}\boxempty P_{3}$
This implies $13+\frac{f_{3}}{4}\geq|F|$. Since $|F|=14$, this implies that $f_{3}\geq 4$. We will frequently make use of this type of combinatorial restriction in this proof, though we will suppress most of the computations moving forward. Combined with Lemma 3.1, we have $4\leq f_{3}\leq 6$.
@@
In light of Proposition 3.2, it will help to more carefully describe how faces can behave in a $2$-cell embedding of $G$ on the torus. We first introduce some necessary notation. We say that a $K_{4}$ fiber of $G$ is an outer $K_{4}$ if its vertices are adjacent to exactly four of the edges contained in the $P_{3}$ fibers. The unique $K_{4}$ fiber of $G$ whose vertices are adjacent to all edges in the $P_{3}$ fibers is called the inner $K_{4}$ fiber. Label the $K_{4}$ fibers $K^{1}$, $K^{2}$, and $K^{3}$, respectively, with $K^{2}$ inner. We use $E_{1}$ to denote the set of edges in the $P_{3}$ fibers incident to both $K^{1}$ and $K^{2}$ and $E_{2}$ to denote the set of edges in the $P_{3}$ fibers incident to both $K^{2}$ and $K^{3}$. Given a $2$-cell embedding of $G$, we say that a quadrilateral is a mixed quadrilateral if it’s boundary is a $4$-cycle of $G$ that consists of two edges from $E_{1}$ ($E_{2}$), one edge from a $K^{1}$ ($K^{3}$), and the corresponding edge from $K^{2}$. Note, that any quadrilateral in an embedding of $G$ that contains a $P_{3}$ fiber edge must be a mixed quadrilateral. Up to symmetry, there are two distinct types of pentagons that bound $5$-cycles in $G$ that could occur in this $2$-cell embedding of $G$, which we denote $p_{1}$ and $p_{2}$. The 5-cycle boundary of $p_{1}$ contains one edge on $K^{2}$, two edges in either $E_{1}$ or $E_{2}$, and two edges in an outer $K_{4}$. Similarly the 5-cycle boundary of $p_{2}$ contains two edges on $K^{2}$, two edges in either $E_{1}$ or $E_{2}$, and one edge in an outer $K_{4}$. Hexagons that bound $6$-cycles in $G$ could possibly be placed in four distinct ways in a $2$-cell embedding of $G$, denoted $h_{1}$, $h_{2}$, $h_{3}$, and $h_{4}$. The 6-cycle boundary of $h_{1}$ contains one edge in $K^{1}$, two edges in $E_{1}$, two edges in $E_{2}$, and one edge in $K^{3}$. The 6-cycle boundary of $h_{2}$ contains two edges in $K^{2}$, two edges in either $E_{1}$ or $E_{2}$, and two edges in an outer $K_{4}$. The 6-cycle boundary of $h_{3}$ contains one edge in $K^{2}$, two edges in either $E_{1}$ or $E_{2}$, and three edges in an outer $K_{4}$. The $6$-cycle boundary of $h_{4}$ contains one edge in an outer $K_{4}$, two edges in either $E_{1}$ or $E_{2}$, and three edges in an outer $K^{2}$.
@@
Lemma 3.3.
If $G$ has a $2$-cell embedding on the torus, then any $n$-gon face of this embedding must be bound by an $n$-cycle of $G$.
@@
Suppose we are given a $2$-cell embedding of $G$ on the torus. Let $F$ be a face in this embedding. Then the boundary of $F$ is a closed walk in $G$ determined by a “facial walk” of this embedding. From Proposition 3.2, we only need to consider when $F$ is a triangle, quadrilateral, pentagon, or hexagon. We will show that if the facial walk for $F$ isn’t an appropriate cycle, then we always reach a contradiction.
First, suppose our embedding admits an $n$-gon face, $F$, that is not bound by an $n$-cycle for some $n=3$, 44, or $5$. The boundary of $F$ is must be a closed walk in $G$, and since the girth of $G$ is three, this closed walk must have a subwalk of the form eveeve for some $e\in E(G)$ and $v\in V(G)$. However, this implies that $G$ admits a degree one vertex, which is a contradiction.
Now, suppose $F$ is a hexagon and the facial walk of $F$ determines a closed walk of length six in $G$ that isn’t a $6$-cycle. If the facial walk for $F$ doesn’t contain any cycle subgraphs, then we once again get a degree one vertex in $G$, which is a contradiction. However, in any $K_{4}$ fiber of $G$, there exist pairs of $3$-cycles that intersect along a single edge. This is the only scenario where the facial walk of $F$ could contain cycle subgraphs and not be a $6$-cycle. However, such an $F$ is not possible. If $f_{6}\geq 1$, then Proposition 3.2 shows that $f_{3}=6$ and $f_{4}=7$. Note that, any triangle in this embedding can only bound edges in $K_{4}$ fibers, and we just assumed our one hexagon only bounds edges in $K_{4}$ fibers. There are only seven quadrilaterals left in this embedding and they must be mixed quadrilaterals, each of which meets exactly two $P_{3}$ fiber edges. However, there are eight edges in our $P_{3}$ fibers, each of which must lie on the boundary of two faces in this embedding, and so, some edge in our $P_{3}$ fiber will only be border one face, which is a contradiction. ∎
@@
We will first show that $G=K_{4}\boxempty P_{3}$ does not embed on the torus by analyzing the four possibilities given in Proposition 3.2. We then show that $G$ does admit an embedding on $S_{2}$.
@@
For sake of contradiction, suppose $G$ 22-cell embeds on the torus with $|F|=14$. In each case, we show that our assumpotions violate Fact (2), giving a necessary contradiction. The following is an important fact we will use throughout the proof: any $e\in E_{1}\cup E_{2}$ does not lie on a $3$-cycle and if such an edge $e$ is bound by a quadrilateral in this embedding, then it must be a mixed quadrilateral.
@@
Suppose $f_{3}=4$ and $f_{4}=10$. To make sure each edge in $E_{1}\cup E_{2}$ is incident to two faces, this embedding must have eight mixed quadrilaterals, four of which are incident to edges in $E_{1}$ and four of which are incident to edges in $E_{2}$. Each of these eight mixed quadrilaterals is incident to exactly one edge in $K^{2}$. Since $12=2|E(K^{2})|$ and each edge must border exactly two faces, four of the (not necessarily distinct) edges in $K^{2}$ must be bound by faces in this embedding that are not mixed quadrilaterals. Since each edge in $E_{1}\cup E_{2}$ already bounds two faces, there must exist a $4$-cycle in $K^{2}$ that is the boundary of a quadrilateral in this embedding. At this point, all faces incident to edges in $K^{2}$ have been determined. Thus, by Lemma 3.1 and our assumptions on $f_{3}$ and $f_{4}$, there are two triangles incident to only edges in $K^{1}$, two triangles incident to only edges in $K^{3}$, and one quadrilateral that must be incident to either only edges in $K^{1}$ or $K^{3}$. Note that $K^{1}$ and $K^{3}$ each have four edges incident to mixed quadrilaterals. Thus, either some edges in $K^{1}$ don’t border two faces and some edges in $K^{2}$ border at least three faces, or vice versa, giving a contradiction.
@@
Suppose $|F|=14$ and $f_{3}=5$, $f_{4}=8$, and $f_{5}=1$. The two types of pentagons ($p_{1}$ and $p_{2}$ described earlier in this section) each bound two edges coming from either $E_{1}$ or $E_{2}$. As a result, we must then have seven mixed quadrilaterals in this embedding to make sure each edge in $E_{1}\cup E_{2}$ borders exactly two faces. This leaves one quadrilateral and five triangles to complete the embedding. Since each mixed quadrilateral is incident to an edge in $K^{2}$ and a pentagon is incident to either one or two edges in $K^{2}$, either four or three more (not necessarily distinct) edges in $K^{2}$ must be incident to the other faces in this embedding. By Lemma 3.1 and $f_{3}=5$, either one or two triangles are incident to only edges in $K^{2}$. Having two triangles contained in $K^{2}$ results in some edge in $K^{2}$ having at least three faces incident to it, which is a contradiction. Having only one triangle contained in $K^{2}$ results in some edge in $K^{2}$ having only one face incident to it, which is a contradiction. The only other possibility is that one triangle and the one (non-mixed) quadrilateral are incident to edges in $K^{2}$, but this would again result in some edge of $K^{2}$ having three faces incident to it. Thus, an embedding of $G$ with $f_{3}=5$, $f_{4}=8$, and $f_{5}=1$ is not possible.

### v3

@@
Abstract
In this paper, we show that if $G$ is $3$-connected, then the Cartesian product of graphs $G\boxempty H$ embeds on the torus if and only if $G$ is outer-cylindrical and $H$ is a path on two vertices, $P_{2}$. As a by-product of our work, we also show that $K_{4}\boxempty P_{3}$ has genus two.
1 Introduction
@@
Classifying embeddings of graphs on surfaces is a well-studied topic in topological graph theory. Here, we are interested in embeddings of a Cartesian product of two graphs $G$ and $H$, denoted $G\boxempty H$, which is defined to have vertex set $V(G)\times V(H)$ and a pair of vertices $(u,v),(u^{\prime},v^{\prime})\in V(G)\times V(H)$ determine an edge in $G\boxempty H$ if and only if $u=u^{\prime}$ and $vv^{\prime}\in E(H)$ or $uu^{\prime}\in E(G)$ and $v=v^{\prime}$. We refer to $G$ and $H$ as the factors of $G\boxempty H$. Every $G\times\{h\}$ with $h\in V(H)$ is called a $G$-fiber of $G\boxempty H$, and $H$-fibers are defined analogously. We refer the reader to ImKlRa2008 for further background on graph Cartesian products. The fiber structures of graph Cartesian products can sometimes be leveraged to efficiently build embeddings and determine genera of such graphs; see MiSa2022 , MoPiWh1990 , Pi1980 , Pi1982 , Pi1989 , Sun2023 , Wh1970 , Wh1971 for some examples. In particular, it would be interesting to know if one can determine the set of all Cartesian products of graphs that embed on a fixed surface. Such a classification was established for planar graphs, that is, graphs that embed in the plane, in the work of Behzad and Mahmoodian BeMa1969 . More recently, the second author along with Abell and McDermott have provided a complete classification of Cartesian products of graphs that embed in the projective plane in AbMcMi2025 .
@@
In this paper, we are interested in classifying which Cartesian product of graphs are toroidal, i.e., embed on the torus. While the classifications of planar and projective planar Cartesian products were reasonably straightforward, we expect the classification of toroidal Cartesian products to be far more complicated. To start, the work of Kuratowski Ku1930 and Wagner Wa1937 showed there are only two minor-minimal obstructions to planarity, namely $K_{5}$ and $K_{3,3}$, and Archdeacon in Ar1980 showed that there is a complete list of 35 graphs that are minor-minimal obstructions to embedding on the projective plane. However, the list of minor-minimal obstructions for embedding on the torus is not completely known, and this list contains at least 17 thousand forbidden minors; see MyWo2018 . In addition, when restricting to Cartesian products of graphs, we see far more flexibility in toroidal embeddings than planar or projective planar embeddings with regards to (vertex) connectivity. If $G\boxempty H$ is planar or projective planar, then both $G$ and $H$ are each at most $2$-connected. This follows from the fact that $K_{4}\boxempty P_{2}$ is both non-planar and non-projective planar, and any $3$-connected graph contains a $K_{4}$ minor; see Lemma 3. However, Example 1 in Section 5 shows that there exist $G\boxempty P_{2}$ that are toroidal where $G$ is $4$-connected. Thus, even a partial classification of toroidal Cartesian products could be a far more complicated task than the planar and projective planar cases.
Here, we provide some major first steps towards this classification problem, with the additional assumption that one factor is $3$-connected. In what follows, a graph $G$ is outer-cylindrical if there exists a planar embedding of $G$ such that the vertices of $G$ all lie on the boundary of at most two faces in this embedding.
Theorem 1.1.
Let $H$ and $G$ be simple, connected, nontrivial graphs, and suppose $G$ is $3$-connected. Then $G\boxempty H$ is toroidal if and only if $G$ is an outer-cylindrical graph and $H=P_{2}$.
The proof of Theorem 1.1 is broken up into a few pieces. In Proposition 5, we show that if $G$ is $3$-connected, then any toroidal $G\boxempty H$ must have $H=P_{2}$. This proof follows from the fact that $K_{4}\boxempty P_{3}$ is not toroidal, which is proved in Theorem 3.1. Proposition 6 provides an explicit construction for building a toroidal embedding of $G\boxempty P_{2}$, where $G$ is outer-cylindrical. Then Proposition 7 shows that if $G\boxempty P_{2}$ is toroidal, then $G$ must be an outer-cylindrical graph. This proof only requires some basic combinatorial topology of graph embeddings.
@@
A significant by-product of our work occurs in Section 3 where we prove that the genus of $K_{4}\boxempty P_{3}$ is two. This work requires a careful combinatorial analysis of all potential embeddings of this graph on a torus, which employs techniques similar to BrSq1988 . In particular, even though the girth of $K_{4}\boxempty P_{3}$ is three, one cannot hope for a triangular embedding since every edge in a $P_{3}$ fiber does not lie on a $3$-cycle. Thus, determining the genus of such a graph is not a straightforward task.
2 Background
@@
In this paper, we assume all graphs are simple, connected, and non-trivial (in the sense of being nonempty and not a single vertex), unless otherwise noted. We use $S_{g}$ to denote an orientable, closed, connected surface of genus $g\in\mathbb{N}\cup\{0\}$. Given an embedding of a graph $G$ on a surface $S_{g}$ via a continuous injective function $f:|G|\rightarrow S_{g}$ where $|G|$ represents a geometric realization of $G$, we sometimes use $f(G)$ to denote this embedding or just $G$ when the embedding is clear from context. If each component of $S_{g}\setminus f(G)$ is homeomorphic to an open disk, then we say this embedding is a $2$-cell embedding. The genus of a graph $G$, denoted $\gamma(G)$, is the minimum $g\in\mathbb{N}\cup\{0\}$ such that $G$ embeds on $S_{g}$. An embedding of $G$ that realizes the minimal genus is called a minimal embedding. We will mainly be interested in $2$-cell embeddings of graphs, and since the work of Youngs Yo1963 shows that each minimal embedding of $G$ is a $2$-cell embedding, we usually do not need to make a distinction when determining the genera of a graph. We refer the reader to GrTu2001 for further background on graph embeddings on surfaces.
@@
Fact 1: Each vertex of $G$ has an open $2$-cell neighborhood in $S_{g}$.
Fact 2: Each edge of $G$ is either incident to two distinct faces or twice incident to one face.
@@
Facts 1 and 2 are general facts about $2$-cell embeddings. Fact 3 follows from Fact 2 and since any edge in a $P_{n}$ fiber is not contained in a $3$-cycle. Fact 4 follows from Fact 2. This collection of facts will play an important role in ruling out potential toroidal embeddings of graphs in Section 3.
A graph $G^{\prime}$ is a minor of $G$ if $G^{\prime}$ can be obtained by a series of vertex deletions, edge deletions, or edge contractions on $G$. Another way to find a lower bound for the genus of a graph $G$ is by finding the genus (or lower bound on the genus) for a subgraph or minor of $G$. Since we will be interested in Cartesian products of graphs in this paper, we first confirm that the product of minors from each factor results in a minor of the product.
Proposition 2.
@@
Let $G$ and $H$ be graphs and suppose $G^{\prime}$ be a minor of $G$. Recall that the vertex sets and edge sets of $G\boxempty H$ are defined as $V(G\boxempty H)=V(G)\times V(H)$ and $E(G\boxempty H)=\{\{(u,v),(u,v^{\prime})\}:u\in V(G),vv^{\prime}\in E(H)\}\cup\{\{(u,v),(u^{\prime},v)\}:uu^{\prime}\in E(G),v\in V(H)\}$. It suffices to show $G^{\prime}\boxempty H$ is a minor of $G\boxempty H$ under an edge deletion, vertex deletion, or edge contraction, which we now handle as three separate cases.
Case 1: Let $G^{\prime}=G-e$ for some $e=uu^{\prime}\in E(G)$. Then $G^{\prime}\boxempty H$ can be obtained from $G\boxempty H$ by deleting edge $\{(u,v),(u^{\prime},v)\}\in E(G\boxempty H)$, for each $v\in V(H)$.
@@
Case 3: Given edge $e=uu^{\prime}\in E(G)$, we contract the edge $e$ in $G$ to create $G^{\prime}$ which now has a new vertex $a\in V(G^{\prime})$ to which $u$ and $u^{\prime}$ have been identified. Then $G^{\prime}\boxempty H$ can be obtained from $G\boxempty H$ by contracting edge $\{(u,v),(u^{\prime},v)\}\in E(G\boxempty H)$, for each $v\in V(H)$. ∎
Corollary 1.
@@
3 The genus of $K_{4}\boxempty P_{3}$
In this section, we will show that $\gamma(K_{4}\boxempty P_{3})=2$. Since every $3$-connected graph contains a $K_{4}$ minor (see Lemma 3), this will play a pivotal role in classifying toroidal Cartesian products where one of the factors is $3$-connected.
@@ 3 The genus of $K_{4}\boxempty P_{3}$
This implies $13+\frac{f_{3}}{4}\geq|F|$. Since $|F|=14$, this implies that $f_{3}\geq 4$. We will frequently make use of this type of combinatorial restriction in this proof, though we will suppress most of the computations moving forward. Combined with Lemma 1, we have $4\leq f_{3}\leq 6$.
@@
In light of Proposition 3, it will help to more carefully describe how faces can behave in a $2$-cell embedding of $G$ on the torus. We first introduce some necessary notation. We say that a $K_{4}$ fiber of $G$ is an outer $K_{4}$ if its vertices are adjacent to exactly four of the edges contained in the $P_{3}$ fibers. The unique $K_{4}$ fiber of $G$ whose vertices are adjacent to all edges in the $P_{3}$ fibers is called the inner $K_{4}$ fiber. Label the $K_{4}$ fibers $K^{1}$, $K^{2}$, and $K^{3}$, respectively, with $K^{2}$ inner. We use $E_{1}$ to denote the set of edges in the $P_{3}$ fibers incident to both $K^{1}$ and $K^{2}$ and $E_{2}$ to denote the set of edges in the $P_{3}$ fibers incident to both $K^{2}$ and $K^{3}$. Given a $2$-cell embedding of $G$, we say that a quadrilateral is a mixed quadrilateral if its boundary is a $4$-cycle of $G$ that consists of two edges from $E_{1}$ ($E_{2}$), one edge from a $K^{1}$ ($K^{3}$), and the corresponding edge from $K^{2}$. Note, that any quadrilateral in an embedding of $G$ that contains a $P_{3}$ fiber edge must be a mixed quadrilateral. Up to symmetry, there are two distinct types of pentagons that bound $5$-cycles in $G$ that could occur in this $2$-cell embedding of $G$, which we denote $p_{1}$ and $p_{2}$. The 5-cycle boundary of $p_{1}$ contains one edge on $K^{2}$, two edges in either $E_{1}$ or $E_{2}$, and two edges in an outer $K_{4}$. Similarly the 5-cycle boundary of $p_{2}$ contains two edges on $K^{2}$, two edges in either $E_{1}$ or $E_{2}$, and one edge in an outer $K_{4}$. Hexagons that bound $6$-cycles in $G$ could possibly be placed in four distinct ways in a $2$-cell embedding of $G$, denoted $h_{1}$, $h_{2}$, $h_{3}$, and $h_{4}$. The 6-cycle boundary of $h_{1}$ contains one edge in $K^{1}$, two edges in $E_{1}$, two edges in $E_{2}$, and one edge in $K^{3}$. The 6-cycle boundary of $h_{2}$ contains two edges in $K^{2}$, two edges in either $E_{1}$ or $E_{2}$, and two edges in an outer $K_{4}$. The 6-cycle boundary of $h_{3}$ contains one edge in $K^{2}$, two edges in either $E_{1}$ or $E_{2}$, and three edges in an outer $K_{4}$. The $6$-cycle boundary of $h_{4}$ contains one edge in an outer $K_{4}$, two edges in either $E_{1}$ or $E_{2}$, and three edges in $K^{2}$.
@@
Lemma 2.
If $G$ admits a $2$-cell embedding on the torus, then any $n$-gon face of this embedding must be bound by an $n$-cycle of $G$.
@@
Suppose we are given a $2$-cell embedding of $G$ on the torus. Let $F$ be a face in this embedding. Then the boundary of $F$ is a closed walk in $G$ determined by a “facial walk” of this embedding. From Proposition 3, we only need to consider when $F$ is a triangle, quadrilateral, pentagon, or hexagon. We will show that if the facial walk for $F$ is not an appropriate cycle, then we always reach a contradiction.
First, suppose our embedding admits an $n$-gon face, $F$, that is not bound by an $n$-cycle for some $n=3$, 44, or $5$. The boundary of $F$ must be a closed walk in $G$, and since the girth of $G$ is three, this closed walk must have a subwalk of the form eveeve for some $e\in E(G)$ and $v\in V(G)$. However, this implies that $G$ admits a degree one vertex, which is a contradiction.
Now, suppose $F$ is a hexagon and the facial walk of $F$ determines a closed walk of length six in $G$ that is not a $6$-cycle. If the facial walk for $F$ does not contain any cycle subgraphs, then we once again get a degree one vertex in $G$, which is a contradiction. However, in any $K_{4}$ fiber of $G$, there exist pairs of $3$-cycles that intersect along a single edge. This is the only scenario where the facial walk of $F$ could contain cycle subgraphs and not be a $6$-cycle. However, such an $F$ is not possible. If $f_{6}\geq 1$, then Proposition 3 shows that $f_{3}=6$ and $f_{4}=7$. Note that, any triangle in this embedding can only bound edges in $K_{4}$ fibers, and we just assumed our one hexagon only bounds edges in $K_{4}$ fibers. There are only seven quadrilaterals left in this embedding and they must be mixed quadrilaterals, each of which meets exactly two $P_{3}$ fiber edges. However, there are eight edges in our $P_{3}$ fibers, each of which must lie on the boundary of two faces in this embedding, and so, some edge in our $P_{3}$ fiber will only be border one face, which is a contradiction. ∎
@@
We will first show that $G=K_{4}\boxempty P_{3}$ does not embed on the torus by analyzing the four possibilities given in Proposition 3. We then show that $G$ does admit an embedding on $S_{2}$.
@@
Suppose $G$ admits a $2$-cell embedding on the torus with $|F|=14$. In each case, we show that our assumptions violate Fact (2), giving a necessary contradiction. The following is an important fact we will use throughout the proof: any $e\in E_{1}\cup E_{2}$ does not lie on a $3$-cycle and if such an edge $e$ is bound by a quadrilateral in this embedding, then it must be a mixed quadrilateral.
@@
Suppose $f_{3}=4$ and $f_{4}=10$. To make sure each edge in $E_{1}\cup E_{2}$ is incident to two faces, this embedding must have eight mixed quadrilaterals, four of which are incident to edges in $E_{1}$ and four of which are incident to edges in $E_{2}$. Each of these eight mixed quadrilaterals is incident to exactly one edge in $K^{2}$. Since $12=2|E(K^{2})|$ and each edge must border exactly two faces, four of the (not necessarily distinct) edges in $K^{2}$ must be bound by faces in this embedding that are not mixed quadrilaterals. Since each edge in $E_{1}\cup E_{2}$ already bounds two faces, there must exist a $4$-cycle in $K^{2}$ that is the boundary of a quadrilateral in this embedding. At this point, all faces incident to edges in $K^{2}$ have been determined. Thus, by Lemma 1 and our assumptions on $f_{3}$ and $f_{4}$, there are two triangles incident to only edges in $K^{1}$, two triangles incident to only edges in $K^{3}$, and one quadrilateral that must be incident to either only edges in $K^{1}$ or $K^{3}$. Note that $K^{1}$ and $K^{3}$ each have four edges incident to mixed quadrilaterals. Thus, either some edges in $K^{1}$ do not border two faces and some edges in $K^{2}$ border at least three faces, or vice versa, giving a contradiction.
@@
Suppose $f_{3}=5$, $f_{4}=8$, and $f_{5}=1$. The two types of pentagons ($p_{1}$ and $p_{2}$ described earlier in this section) each bound two edges coming from either $E_{1}$ or $E_{2}$. As a result, we must then have seven mixed quadrilaterals in this embedding to make sure each edge in $E_{1}\cup E_{2}$ borders exactly two faces. This leaves one quadrilateral and five triangles to complete the embedding. Since each mixed quadrilateral is incident to an edge in $K^{2}$ and a pentagon is incident to either one or two edges in $K^{2}$, either four or three more (not necessarily distinct) edges in $K^{2}$ must be incident to the other faces in this embedding. By Lemma 1 and $f_{3}=5$, either one or two triangles are incident to only edges in $K^{2}$. Having two triangles bordered by edges in $K^{2}$ results in some edge in $K^{2}$ having at least three faces incident to it, which is a contradiction. So, suppose there is only one triangle bordered by edges in $K^{2}$, and so, each outer $K_{4}$ is bordered by two triangles. In this case, a $p_{2}$ pentagon must occur to make sure each edge of $K^{2}$ is bordered by exactly two faces. This also means the one remaining quadrilateral must be incident to only edges in a single outer $K_{4}$, say $K^{1}$. Then some edge in $K^{1}$ has three faces incident to it. Thus, an embedding of $G$ with $f_{3}=5$, $f_{4}=8$, and $f_{5}=1$ is not possible.